the potential difference across the terminals of a battery of emf 12 V and internal resistance 2 Ω drops to 10 V when it is connected to a silver voltmeter. Find the silver deposited at the cathode in half an hour. Atomic weight of silver is 107.9 g/mole.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. For potential drop, t = 30 min = 180 sec V base i = V base f + iR ⇒ 12 = 10 + 2i ⇒ i = 1 Amp m = Zit = 107.9/96500 * 1*30*60 = 201 g = 2 g